# BEV-LIO(LC): BEV Image Assisted LiDAR-Inertial Odometry with Loop Closure

### v1

@@ Abstract
This work introduces BEV-LIO(LC), a novel LiDAR-Inertial Odometry (LIO) framework that combines Bird’s Eye View (BEV) image representations of LiDAR data with geometry-based point cloud registration and incorporates loop closure (LC) through BEV image features. By normalizing point density, we project LiDAR point clouds into BEV images, thereby enabling efficient feature extraction and matching. A lightweight convolutional neural network (CNN) based feature extractor is employed to extract distinctive local and global descriptors from the BEV images. Local descriptors are used to match BEV images with FAST keypoints for reprojection error construction, while global descriptors facilitate loop closure detection. Reprojection error minimization is then integrated with point-to-plane registration within an iterated Extended Kalman Filter (iEKF). In the back-end, global descriptors are used to create a KD-tree-indexed keyframe database for accurate loop closure detection. When a loop closure is detected, Random Sample Consensus (RANSAC) computes a coarse transform from BEV image matching, which serves as the initial estimate for Iterative Closest Point (ICP). The refined transform is subsequently incorporated into a factor graph along with odometry factors, improving the global consistency of localization. Extensive experiments conducted in various scenarios with different LiDAR types demonstrate that BEV-LIO(LC) outperforms state-of-the-art methods, achieving competitive localization accuracy. Our code, video and supplementary materials can be found at https://github.com/HxCa1/BEV-LIO-LC.
@@ III-A Symbol Definition
We denote the world frame as , the LiDAR frame as and the IMU frame as . Transformation from LiDAR frame to IMU frame is represented as . The robot state in -th LiDAR frame is defined as , with and representing the biases of gyroscope and accelerometer.
@@ III-E Reprojection Residual & Kalman Update
where denotes and is a factor that scales the projection transformation to maintain consistency across different directions.
@@ III-F Loop Closure
Our back-end adopts a factor graph framework similar to LIO-SAM [12], integrating loop closure constraints through the following pipeline:
@@ IV-B Loop Closure Evaluation
In summary, our method demonstrates superior performance compared to direct point cloud matching methods like [16] and spherical projection methods such as [29]. While these methods perform well in certain conditions, BEV-LIO(LC) achieves higher localization precision by coupling BEV image feature and a back-end optimization. This highlights the advantage of our method in improving localization accuracy and loop closure performance for correcting accumulated errors.
@@ V Conclusion
In this work, we propose BEV-LIO(LC), a LIO framework that combines BEV image representations of LiDAR data with geometric registration for front-end odometry and integrates loop closure via BEV image features and factor graph optimization for the back-end. Our approach enhances front-end odometry by combining reprojection error from BEV image matching with geometric registration, and utilizes iEKF to couple these residuals, significantly improving odometry accuracy and reliability. For loop closure, our back-end system employs a KD-tree-indexed BEV descriptor database for efficient detection. Upon loop detection, a coarse transform computed via RANSAC from BEV image matching initializes the ICP process, which refines the transform for improved global consistency of localization when integrated into a factor graph alongside odometry factors. Extensive experiments across various scenarios and with different LiDAR types validate the competitive localization accuracy of BEV-LIO(LC) compared to state-of-the-art methods. Looking ahead, future work will focus on eliminating the dependency on CNNs for feature extraction, with the goal of further enhancing real-time performance of the system.

### v2

@@ Abstract
This work introduces BEV-LIO(LC), a novel LiDAR-Inertial Odometry (LIO) framework that combines Bird’s Eye View (BEV) image representations of LiDAR data with geometry-based point cloud registration and incorporates loop closure (LC) through BEV image features. By normalizing point density, we project LiDAR point clouds into BEV images, thereby enabling efficient feature extraction and matching. A lightweight convolutional neural network (CNN) based feature extractor is employed to extract distinctive local and global descriptors from the BEV images. Local descriptors are used to match BEV images with FAST keypoints for reprojection error construction, while global descriptors facilitate loop closure detection. Reprojection error minimization is then integrated with point-to-plane registration within an iterated Extended Kalman Filter (iEKF). In the back-end, global descriptors are used to create a KD-tree-indexed keyframe database for accurate loop closure detection. When a loop closure is detected, Random Sample Consensus (RANSAC) computes a coarse transform from BEV image matching, which serves as the initial estimate for Iterative Closest Point (ICP). The refined transform is subsequently incorporated into a factor graph along with odometry factors, improving the global consistency of localization. Extensive experiments conducted in various scenarios with different LiDAR types demonstrate that BEV-LIO(LC) outperforms state-of-the-art methods, achieving competitive localization accuracy. Our code and video can be found at https://github.com/HxCa1/BEV-LIO-LC.
@@ III-A Symbol Definition
We denote the world frame as , the LiDAR frame as and the IMU frame as . Transformation from LiDAR frame to IMU frame is represented as . The robot state at the -th LiDAR frame is defined as , where denotes the orientation, represents the position, is the linear velocity, and indicate the gyroscope and accelerometer biases, respectively, and denotes the gravity vector.
@@ III-E Reprojection Residual & Kalman Update
where is a factor that scales the projection transformation to maintain consistency across different directions.
@@ III-F Loop Closure
Our back-end adopts a factor graph framework FAST-LIO-SAM[36], integrating loop closure constraints through the following pipeline:
@@ IV-B Loop Closure Evaluation
In summary, our method demonstrates superior performance compared to direct point cloud matching methods like [16] and spherical projection methods like [29]. While they perform well in certain conditions, BEV-LIO(LC) achieves higher localization precision by coupling BEV image feature and a back-end optimization. This highlights the advantage of our method in improving localization accuracy and loop closure performance for correcting accumulated errors.
@@ V Conclusion
In this work, we propose BEV-LIO(LC), a LIO framework that combines BEV images of LiDAR data with geometric registration for front-end odometry and integrates loop closure via BEV image features and factor graph optimization for the back-end. Our approach enhances front-end odometry by combining reprojection error from BEV image matching with geometric registration, and utilizes iEKF to couple these residuals, improving odometry accuracy and reliability. For loop closure, our back-end system employs a KD-tree-indexed BEV descriptor database. Upon loop detection, a coarse transform computed via RANSAC from BEV image matching initializes the ICP process, which refines the transform for improved global consistency of localization when integrated into a factor graph alongside odometry factors. Extensive experiments across various scenarios and with different LiDAR validate the competitive localization accuracy of BEV-LIO(LC) compared to state-of-the-art methods. Looking ahead, future work will focus on eliminating the dependency on CNNs for feature extraction, with the goal of further enhancing real-time performance of the system.